# Relativistic distorted-wave analysis of the missing-energy spectrum measured with monochromatic $\nu_{\mu}$12C interactions at JSNS2

### 03 Mar 2026 v1

@@ V Comparison with experimental data
The comparison between the results shown in the two panels clarifies the role of the residual-nucleus recoil. As already pointed out, neglecting this term in the evaluation of $E_{\rm vis}$ (see Eq. (18)) leads to an overestimation of the missing energy. In our calculations, this corresponds to a shift of approximately 1 to 5 MeV. As a consequence, when the recoil is not included, the p-shell peak appears in the $20<E_{m}<25$ MeV bin, as shown in the lower panel. If the recoiling residual nucleus does not produce visible light in the detector, this effect could help explain the discrepancies observed between the data and the generator predictions, which does not account for recoil effects in the kinematics, in the region of the peak [3].
@@
The relevance of neutron emission is clearly illustrated in Fig. 5 by comparing the ED-RMF + NuWro results with those of the other calculations. The most prominent effect is the migration of events from the region dominated by the p-shells ($E_{m}<22$ MeV) toward higher values of $E_{m}$, although the other shells are also affected by this mechanism. As a result, the ED-RMF + NuWro predictions overestimate the experimental data for $E_{m}>30$ MeV in both cases presented in the figure, namely with and without including the recoil of the residual nucleus.
@@
All the theoretical cross sections shown in Fig. 5 vanish below the neutron emission threshold of 18.7 MeV, as expected within our RDWIA models. However, the experimental data exhibit nonzero values even at energies below this threshold.
@@
The results obtained with this smearing procedure are shown in Fig. 6. We observe that the region below threshold is now populated by the theoretical cross sections, and the agreement with the experimental data is correspondingly improved.
@@ VI Summary and conclusions
The situation is quite different for the ED-RMF + NuWro results. The shape of the distribution is modified, as part of the strength is shifted from the main $p$-shell contribution to larger missing-energy values. This effect arises from neutron emission induced by the re-scattering of the primary proton. Since the neutron energy is not detected by the JSNS2 detector, the visible energy $E_{\rm vis}$ is reduced, consequently leading to a larger reconstructed missing energy (see Eqs. (17) and (18)). Including this mechanism leads to poorer agreement with the data, unlike the ED-RMF result. Ref. [34] compares proton–12C reaction cross-section with the predictions of different neutrino event generators and finds sizable discrepancies among the generators for $T_{p}<30$ MeV, the kinematic region relevant for JSNS2. Nevertheless, even when a generator reproduces the low-$T_{p}$ reaction cross-section, the observable is inclusive and does not ensure an accurate modeling of particular inelastic final states, such as channels involving neutron emission.
@@ Acknowledgements.
This work has been partially supported by INFN under Project NUCSYS. We would like to thank J. García-Marcos and R. González-Jiménez for propagating our theoretical predictions through NuWro and for useful discussions and V. Belocchi for providing the MEC calculation.

### 05 Mar 2026 v2

@@ V Comparison with experimental data
The comparison between the results shown in the two panels clarifies the role of the residual-nucleus recoil. As already pointed out, neglecting this term in the evaluation of $E_{\rm vis}$ (see Eq. (18)) leads to an overestimation of the missing energy. In our calculations, this corresponds to a shift of approximately 1 to 5 MeV. As a consequence, when the recoil is not included, a large part of the p-shell strength appears in the $20<E_{m}<25$ MeV bin, as shown in the lower panel.
@@
The relevance of neutron emission is clearly illustrated in Fig. 5 by comparing the ED-RMF + NuWro results with those of the other calculations. The most prominent effect is the migration of events from the region dominated by p-shells ($E_{m}<22$ MeV) to higher values of $E_{m}$, although the other shells are also affected by this mechanism. As a result, the ED-RMF + NuWro predictions overestimate the experimental data for $E_{m}>30$ MeV in both cases presented in the figure, namely with and without including the recoil of the residual nucleus.
@@
All theoretical cross sections shown in Fig. 5 vanish below the neutron emission threshold of 18.7 MeV, as expected within our RDWIA models. However, the experimental data exhibit nonzero values even at energies below this threshold.
@@
The results obtained with this smearing procedure are shown in Fig. 6. We observe that the region below the threshold is now populated by the theoretical cross sections, and the agreement with the experimental data is improved, especially for the case including the recoil. In the lower panel of Fig. 6, the position of the peak is no longer visible around $E_{m}\approx 18.7$ MeV due to the combined effect of neglecting the recoil and the detector smearing. If the recoiling residual nucleus does not produce visible light in the detector, this combined effect could help explain the discrepancies observed between the data and the generator predictions, which does not account for recoil effects in the kinematics, in the region of the peak [3].
@@ VI Summary and conclusions
The situation is quite different for the ED-RMF + NuWro results. The shape of the distribution is modified, as part of the strength is shifted from the main $p$-shell contribution to larger missing-energy values. This effect arises from neutron emission induced by the re-scattering of the primary proton. Since the neutron energy is not detected by the JSNS2 detector, the visible energy $E_{\rm vis}$ is reduced, consequently leading to a larger reconstructed missing energy (see Eqs. (17) and (18)). Including this mechanism leads to poorer agreement with the data, unlike the ED-RMF result. Ref. [34] compares proton–12C reaction cross-section with the predictions of different neutrino event generators and finds sizable discrepancies among the generators for $T_{p}<50$ MeV, the kinematic region relevant for JSNS2. Nevertheless, even when a generator reproduces the low-$T_{p}$ reaction cross-section, the observable is inclusive and does not ensure an accurate modeling of particular inelastic final states, such as channels involving neutron emission.
@@ Acknowledgements.
This work has been partially supported by INFN under Project NUCSYS.